OpenStudy (anonymous):

which of the following accurately depicts the transformation of y=x^2 to the function below? y=5(x-2)^2+1 a. shift up 1 unit, stretch horizontally by a factor of 5, then shift 2 units b. shift right 2 units, stretch vertically by a factor of 5, then shift up 1 unit c. shift 5 units right, stretch vertically by factor of 2, shift up 1 d. shift right 2 units, shrink vertically to 1/5 of the original height, shift up 1

OpenStudy (anonymous):

@RhondaSommer

OpenStudy (jdoe0001):

\(\qquad \qquad \qquad \qquad \textit{function transformations} \\ \quad \\ \begin{array}{rllll} % left side templates f(x)=&{\color{purple}{ A}}({\color{blue}{ B}}x+{\color{red}{ C}})+{\color{green}{ D}} \\ \quad \\ y=&{\color{purple}{ A}}({\color{blue}{ B}}x+{\color{red}{ C}})+{\color{green}{ D}} \\ \quad \\ f(x)=&{\color{purple}{ A}}\sqrt{{\color{blue}{ B}}x+{\color{red}{ C}}}+{\color{green}{ D}} \\ \quad \\ f(x)=&{\color{purple}{ A}}\mathbb{R}^{{\color{blue}{ B}}x+{\color{red}{ C}}}+{\color{green}{ D}} \end{array}\qquad \begin{array}{llll} % right side info \bullet \textit{ stretches or shrinks horizontally by } {\color{purple}{ A}}\cdot {\color{blue}{ B}}\\ \bullet \textit{ horizontal shift by }\frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\\ \qquad if\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\textit{ is negative, to the right}\\ \qquad if\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\textit{ is positive, to the left}\\ \bullet \textit{vertical shift by }{\color{green}{ D}}\\ \qquad if\ {\color{green}{ D}}\textit{ is negative, downwards}\\ \qquad if\ {\color{green}{ D}}\textit{ is positive, upwards} \end{array} \\ \quad \\ \quad \\ \bf y={\color{purple}{ 5}}(x{\color{red}{ -2}})^2{\color{green}{ +1}}\) any ideas?

OpenStudy (anonymous):

is it C?

OpenStudy (jdoe0001):

what makes you think is C?

OpenStudy (anonymous):

no wait is it A because it shift horizontally and shifts up not down?

OpenStudy (anonymous):

and does it shift left because 2 is negative? i think it's A

OpenStudy (jdoe0001):

hmmmm notice the template, what happens when C/B is negative

OpenStudy (anonymous):

ohh it shifts to the right

OpenStudy (jdoe0001):

notice, you have x -2 or 1x-2 meaning B = 1, C= -2

OpenStudy (jdoe0001):

yeap expands by a factor of 5 shift horizontally to the "right" by 2 units and goes up by 1

OpenStudy (anonymous):

so it's D

OpenStudy (jdoe0001):

mmm actually a factor of 5, shrinks anyhow shrinks it by a factor of 5 shift horizontally to the "right" by 2 units and goes up by 1

OpenStudy (jdoe0001):

but check your choices more closely

OpenStudy (anonymous):

the only one that shrinks is D??

OpenStudy (jdoe0001):

hmmmm hmmm I think.... yes.. you're correct.... it'd be D then

OpenStudy (anonymous):

thank you!

OpenStudy (jdoe0001):

yw